OpenStudy (anonymous):

For the followign system of equations, what is the x- value of the solution? -x+2y=6 6y=x+18 And please show your work with this

OpenStudy (anonymous):

-x+2y=6 -x=6-2x x=-6+2y

OpenStudy (anonymous):

then substitute it for x in the next

OpenStudy (anonymous):

6y =(-6+2y)+18

OpenStudy (anonymous):

solve for y 6y= 12+2y 4y=12 y=3

OpenStudy (hunus):



OpenStudy (hunus):

The solution will be where their values intersect

OpenStudy (anonymous):

then with the y value sub it into one. -x+2y=6 -x+2(3)=6 -x+6=6 x=0